OpenStudy (anonymous):

slove. -2(3-m)+15=6m-4(m-20). show all work

OpenStudy (anonymous):

-2(3-m)+15=6m-4(m-20) -(6-2m) +15 = 6m -(4m-80) -6+2m +15 = 6m -4m +80 2m+ 9 = 2m + 80 9=80 Since 9 is not equals to 80. There is no solution for m.